OpenStudy (anonymous):

plzz help me now....... if cosy=xcos(a+y) theprove that dy/dx=(cos^2 (a+y))/sina

OpenStudy (anonymous):

i jst dbtknw ...y i g8 stuck wit simple sums....:(

OpenStudy (dumbcow):

using implicit differentiation: \[\frac{dy}{dx} = \frac{\cos(a+y)}{x \sin(a+y)-\sin(y)}\]

OpenStudy (dumbcow):

then im guessing use sum of angles formula

OpenStudy (dumbcow):

sin(a+y) = sin(a)cos(y) + sin(y)cos(a)

OpenStudy (anonymous):

\[shouldnt ittt be -\sin(y) dy/dx =-\sin(a+y)x+coas(a+y)\]

OpenStudy (dumbcow):

i don't know if im helping :| there should be a dy/dx with the -sin(a+y)x term

OpenStudy (anonymous):

ywesssss it shud b sry it din write

OpenStudy (anonymous):

pellet....yes u r write....mdumb :P

OpenStudy (anonymous):

*pellettt

OpenStudy (anonymous):

*pelletttt

OpenStudy (anonymous):

***crap...

OpenStudy (anonymous):

censorship even here??? :P

OpenStudy (dumbcow):

haha from here though im not sure where to go, they don;t have an x in their answer ...i don't see how the x cancels anywhere